OpenStudy (anonymous):

how do you solve percentages of decimals like the 36% of 4.58?

OpenStudy (anonymous):

a percentage of another number is the same as multiplying that percentage by the number. Like 30% of 10 is the same as 30% * 10... which is like 0.30 * 10 = 3 It is the same idea when finding a percentage of a decimal number... just multiply the percentage times the other number.

OpenStudy (campbell_st):

a percentage is another way of writing a fraction of decimal... you can change your percentage to a fraction 36/100 or a decimal 0.36 and just multiply your quantity 0.36 x 4.58 = or 36/100 x 4.58 an alternate method is to find 10% 10% of 4.58 = 0.458 so 30% = 3 x 0.458 5% = 0.458/2 1% = 0.0458 ------------- add them for 36% =